OpenStudy (anonymous):

One fifth of a hive of bees flew to the daisies, one third to the tulips, and three times the difference of these two numbers flew to the arbor. One bee continued to fly about the rose bush. How many bees were there?

OpenStudy (anonymous):

There were 15 bees in the hive. Let h be the number of bees in the hive. Solve the following for h:\[\frac{h}{5}+\frac{h}{3}+3\left(\frac{h}{3}-\frac{h}{5}\right)+1=h\]h = 15 bees

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

try to solve my age problem while you are at it no one else has gotten it yet

OpenStudy (anonymous):

it's just a bit harder

OpenStudy (anonymous):

Where is the age problem. What is it's web address?

OpenStudy (anonymous):

look at the questions asked on my profile... only been two